OpenStudy (anonymous):

i need biiig help ! i will post question when someone willing is going to help. it looks like a long thing but its really not. the actual question is short. anyone up for the challenge ?

OpenStudy (aaronq):

ill help

OpenStudy (anonymous):

Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? Follow the 5 steps below to complete this problem. Step 1 Pick a friend or family member to be the character of your word problem. This friend or family member may do one of the following: Drive a boat [[Drive a jet ski]] Step 2: Select a current speed of the water in mph. 50 mph Step 3: Select the number of hours that your friend or family member drove the boat or jets ski against the current speed you chose in step 2. 1 hour Step 4: Select the number of hours that your friend or family member made the same trip with the current (this should be a smaller number, as your friend or family member will be traveling with the current). 40 mph Step 5: Write out the word problem you created and calculate how fast your friend or family member was traveling in still water. Round your answer to the nearest mph.

OpenStudy (aaronq):

you should post one at a time, il address the first one. Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? let x be tonys age, and y be belindas age, so their combined ages: (1) x+y=56 and belinda is 8 more than twice his age meaning: (2) y=2x +8 now isolate in equation (1), isolate either variable, i chose x: x=56-y substitute it into (2) y=2(56-y)+8 solve for y (belindas age), then knowing y, find x using any of the previous equations ie. (1) or (2)

OpenStudy (anonymous):

okay i was with ya until the last part. are u asking me to tell u the answer, 1 or 2, or is that the rest of it ? and yeaa sorry im in a hurry and didnt think of that.

OpenStudy (aaronq):

they're asking you for their ages, so y will be belindas age and x will be tonys. solve for y, y=2(56-y)+8 then plug y into (1) x+y=56 or (2) y=2x +8 (so that you only have one unknown) then find x

OpenStudy (anonymous):

@aaronq ????

OpenStudy (aaronq):

they're asking you for their ages, so y will be belindas age and x will be tonys. solve for y, y=2(56-y)+8 then plug y into (1) x+y=56 or (2) y=2x +8 (so that you only have one unknown) then find x

OpenStudy (aaronq):

i set it up for ya, but haven't solved for the answer thats your job

OpenStudy (anonymous):

okay but idk how to plug in y. ill plug it into (1) can u tell me how and ill do it and tell u what i get ?

OpenStudy (aaronq):

solve for y, it will be a number, lets say 4 (this is an example not the answer) then plug it into (1) x+4 =56, find x

OpenStudy (anonymous):

i dont know how. -___- do i multiply ? add ? divide ? what ?

OpenStudy (aaronq):

you don't know how to solve this y=2(56-y)+8?

OpenStudy (anonymous):

not at all. sorry.

OpenStudy (anonymous):

do i use pemdas ?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

i got y = 120

OpenStudy (aaronq):

lol do you think belinda is 120 years old? hint: their ages have to add up to 56, remember? haha y=2(56-y)+8 expand bracket y=112-2y+8 bring 2y to other side 2y+y=112+8 add up y's and number 3y=120 isolate y y=120/3 y=40

OpenStudy (anonymous):

oh im a fail :/

OpenStudy (aaronq):

haha it's okay, you just gotta practise

OpenStudy (anonymous):

i will. so is there more to do ?

OpenStudy (aaronq):

yep, since you know the value of y, you can use it in any of the previous equations to find x. (1) x+y=56 x+40=56 what is x?

OpenStudy (anonymous):

x+40=56

OpenStudy (aaronq):

yes, solve for x